If x + y = 4
what is the perimeter of a square with side length 2x^2 + 4xy + 2y^2?

Answer by jim_thompson5910(35256)


2x^2 + 4xy + 2y^2

2(x^2 + 2xy + y^2)

2(x + y)^2

2(4)^2

2(16)

32

So if x+y = 4, then 2x^2 + 4xy + 2y^2 = 32

So the side length is 32 units.

The perimeter is then 32*4 = 128 units